# Tracking Brownian fluid particles in large eddy simulations

### v2

@@ 5 Appendix: Fluid flows represented by Brownian fluid particles
Given $\eta\in\mathbb{R}^{d}$, the distribution of $X$, a weak solution to (51) that $X_{0}=\eta$, is denoted by $\mathbb{P}^{\eta}$, which is a probability measure on the space $C$([0,)$,\mathbb{R}^{d})$ ) of continuous paths. Consider for every $\xi$ , Itô’s stochastic differential equation
@@
where $B_{t}$ is $d$-dimensional Brownian motion. In addition to Itô’s SDE formulation, we shall use the weak solution formulation as well, cf. [18]. Let $\varOmega=C([0,\infty),\mathbb{R}^{d})$ be the continuous path space in $\mathbb{R}^{d}$, and $X=(X_{t})_{t\geq 0}$ denote the coordinate process, that is, for each $t\geq 0$, $X_{t}:\varOmega\mapsto\mathbb{R}^{d}$, which sends each path $\psi\in\varOmega$ to $\psi(t)$. For simplicity, $X_{t}$ shall be written as $\psi(t)$ if no confusion may arise from the context. Let $\mathcal{F}_{t}^{0}=\sigma\{X_{s}:s\leq t\}$ be the smallest $\sigma$-algebra on $\varOmega$ such that $X_{s}$ are measurable for all $s\leq t$, and $\mathcal{F}^{0}=\sigma\{X_{s}:s<\infty\}$. Then $\mathcal{F}^{0}=\mathcal{B}(\varOmega)$ the Borel $\sigma$-algebra on $\varOmega$ generated by the uniform convergence over any bounded subset of $[0,\infty)$.
@@
$\left(L_{-b}-\frac{\partial}{\partial t}\right)\varPsi+g=0\quad\textrm{ in }\mathbb{R}_{+}^{d}$ (56)
@@
$+\int_{0}^{t}\int_{\mathbb{R}_{+}^{d}}\mathbb{P}_{b}^{\eta,0\rightarrow\xi,t}\left[1_{\{s>\gamma_{t}(X^{\eta})\}}g(\psi(s),s)\right]p_{b}(0,\eta;t,\xi)\textrm{d}\eta\textrm{d}s$ (57)
@@
$\textrm{d}Y_{t}=-b(Y_{t},T-t)\textrm{d}t+\sqrt{2\nu}\textrm{d}B_{t},\quad Y_{0}=\xi$ (58)
@@
$Y_{t\wedge T_{\xi}}=\xi-\int_{0}^{t}1_{\left\{s<T\wedge T_{\xi}\right\}}b(Y_{s},T-s)\textrm{d}s+\sqrt{2\nu}\int_{0}^{t}1_{\left\{s<T\wedge T_{\xi}\right\}}\textrm{d}B_{s}$ (60)
@@
$+1_{\left\{t<T_{\xi}\right\}}\left(\nu\Delta-b\cdot\nabla-\frac{\partial}{\partial t}\right)\varPsi(Y_{t},T-t)\textrm{d}t$ (61)
@@
$=\int_{0}^{T}\int_{\mathbb{R}^{3}}\mathbb{P}_{b}^{\eta,0\rightarrow\xi,T}\left[1_{\{T-t<\zeta(\psi\circ\tau_{T})\}}g(\psi(t),t)\right]p_{b}(0,\eta;T,\xi)\textrm{d}\eta\textrm{d}t$
@@
$+\int_{0}^{T}\int_{\mathbb{R}^{3}}\mathbb{P}_{b}^{\eta,0\rightarrow\xi,T}\left[1_{\{T-t<\zeta(\psi\circ\tau_{T})\}}g(\psi(t),t)\right]p_{b}(0,\eta;T,\xi)\textrm{d}\eta\textrm{d}t$ (64)
@@
$+\int_{0}^{t}\int_{\mathbb{R}^{d}}\mathbb{P}_{b}^{\eta,0\rightarrow\xi,t}\left[f(X_{s},s)\right]p_{b}(0,\eta;t,\xi)\mathrm{d}\eta\mathrm{d}s$ (66)

### v4

@@ Appendix A Fluid flows represented by Brownian fluid particles
[…]
@@
where $B_{t}$ is $d$-dimensional Brownian motion. In addition to Itô’s SDE formulation, we shall use the weak solution formulation as well, cf. [27]. Let $\varOmega=C([0,\infty),\mathbb{R}^{d})$ be the continuous path space in $\mathbb{R}^{d}$, and $X=(X_{t})_{t\geq 0}$ denote the coordinate process, that is, for each $t\geq 0$, $X_{t}:\varOmega\mapsto\mathbb{R}^{d}$, which sends each path $\psi\in\varOmega$ to $\psi(t)$. For simplicity, $X_{t}$ shall be written as $\psi(t)$ if no confusion may arise from the context. Let $\mathcal{F}_{t}^{0}=\sigma\{X_{s}:s\leq t\}$ be the smallest $\sigma$-algebra on $\varOmega$ such that $X_{s}$ are measurable for all $s\leq t$, and $\mathcal{F}^{0}=\sigma\{X_{s}:s<\infty\}$. Then $\mathcal{F}^{0}=\mathcal{B}(\varOmega)$ the Borel $\sigma$-algebra on $\varOmega$ generated by the uniform convergence over any bounded subset of $0,\infty)$.
@@
$\left(L_{-b}-\frac{\partial}{\partial t}\right)\varPsi+g=0\quad\textrm{ in }\mathbb{R}_{+}^{d},$ (64)
@@
$+\int_{0}^{t}\int_{\mathbb{R}_{+}^{d}}\mathbb{P}_{b}^{\eta,0\rightarrow\xi,t}\left[1_{\{s>\gamma_{t}(X^{\eta})\}}g(\psi(s),s)\right]p_{b}(0,\eta;t,\xi)\textrm{d}\eta\textrm{d}s,$ (65)
@@
$\textrm{d}Y_{t}=-b(Y_{t},T-t)\textrm{d}t+\sqrt{2\nu}\textrm{d}B_{t},\quad Y_{0}=\xi,$ (66)
@@
$Y_{t\wedge T_{\xi}}=\xi-\int_{0}^{t}1_{\left\{s<T\wedge T_{\xi}\right\}}b(Y_{s},T-s)\textrm{d}s+\sqrt{2\nu}\int_{0}^{t}1_{\left\{s<T\wedge T_{\xi}\right\}}\textrm{d}B_{s},$ (68)
@@
$+1_{\left\{t<T_{\xi}\right\}}\left(\nu\Delta-b\cdot\nabla-\frac{\partial}{\partial t}\right)\varPsi(Y_{t},T-t)\textrm{d}t,$ (69)
@@
$=\int_{0}^{T}\int_{\mathbb{R}^{3}}\mathbb{P}_{b}^{\eta,0\rightarrow\xi,T}\left[1_{\{T-t<\zeta(\psi\circ\tau_{T})\}}g(\psi(t),t)\right]p_{b}(0,\eta;T,\xi)\textrm{d}\eta\textrm{d}t,$
@@
$+\int_{0}^{T}\int_{\mathbb{R}^{3}}\mathbb{P}_{b}^{\eta,0\rightarrow\xi,T}\left[1_{\{T-t<\zeta(\psi\circ\tau_{T})\}}g(\psi(t),t)\right]p_{b}(0,\eta;T,\xi)\textrm{d}\eta\textrm{d}t.$ (72)
@@
$+\int_{0}^{t}\int_{\mathbb{R}^{d}}\mathbb{P}_{b}^{\eta,0\rightarrow\xi,t}\left[f(X_{s},s)\right]p_{b}(0,\eta;t,\xi)\mathrm{d}\eta\mathrm{d}s,$ (74)